OpenStudy (anonymous):

Please help! Simplify: 10/x+1 ------- 1/2-7/x+1

OpenStudy (anonymous):

hard to read. maybe \[\frac{\frac{10}{x+1}}{\frac{1}{2}-\frac{7}{x+1}}\]?

OpenStudy (anonymous):

2x+20/3x-14

OpenStudy (anonymous):

Yes! :) That's it.

OpenStudy (anonymous):

if so, multiply top and bottom by x + 1 carefully. you get \[\frac{10}{\frac{1}{2}(x+1)-7}\]

OpenStudy (anonymous):

10/x+15? 10/x-13?

OpenStudy (anonymous):

now maybe multiply top and bottom by 2 to get \[\frac{20}{(x+1)-14}\] \[\frac{20}{x-13}\]

OpenStudy (anonymous):

Oh, I see. Thank you so much! :)

OpenStudy (anonymous):

yw